Sale of Equipment Equipment was acquired at the beginning of the year at a cost of $662,500. The equipment was depreciated usplease explain to me the calculations and where i went wrong... thank you
Answer #1

a.

Depreciation as per straight line method = Cost of equipment - Residual value / useful life

Depreciation as per straight line method = $662,500 - 46,080 / 9 = $68,491 per year

Depreciation for first year = $68,491

b.

Accumulated depreciation = Total depreciation for 8 years

Accumulated depreciation at the end of eight year = $68,491 * 8 = $547,928

Book value of equipment at the end of eight year = $662,500 - 547,928 = $114,572

Gain (loss) on sale of equipment at the end of eight year = Sale value - Book value of equipment at the end of eight year

Gain (loss) on sale of equipment at the end of eight year = $107,811 - 114,572 = $(6,761) Loss

c.

Cash $107,811
Accumulated depreciation - Equipment 547,928
Loss on sale of equipment 6,761
Equipment $662,500